OpenStudy (anonymous):

6 2/3% is what as a fraction? (reduce the fraction if possible)

OpenStudy (anonymous):

Do you mean \[6 \frac{2}{3}\]? Or something else?

OpenStudy (anonymous):

yes i do

OpenStudy (anonymous):

...........i think.

OpenStudy (anonymous):

ummm

OpenStudy (anonymous):

%=x/100

OpenStudy (anonymous):

so would it be 6 2/3 then?

OpenStudy (anonymous):

20/3 is the equivalence

hero (hero):

6 2/3 % = 20/3 * (1/100) = 20/300 = 2/30 = 1/15